OpenStudy (anonymous):

Choose the correct simplification of the expression (4x3y2z4)(2x3y4z3). 8x9y8z12 6x6y6z7 8x6y6z7 6x9y8z12

OpenStudy (anonymous):

Multiply the coefficients first, then use the law of indice, a^m • a^n = a^(m + n)

OpenStudy (anonymous):

Order and grouping doesn't matter with multiplication so you can re-write the expression to look like this: \(4x^3 \times y^2\times z^4\times2x^3 \times y^4 \times z^3\) Assuming I have understood your original question.... Please use ^ to indicate an exponent \(x^2\) can be written as x^2